OpenStudy (anonymous):

i dont know what to do |10x^4y^2|^3 over |-5x^3y^2|

OpenStudy (pfenn1):

\[|10x^4y^2|^3 \over |-5x^3y^2|\] Do you know how to simplify this?

OpenStudy (anonymous):

Distribute the exponent, then simplify the like terms:

OpenStudy (anonymous):

@davisgurl101 Can you follow our instruction?

OpenStudy (anonymous):

no i do not i just started this class to get my GPA up and idk how to do it

OpenStudy (anonymous):

and yes i can

OpenStudy (anonymous):

(x^2)^3 = ...

OpenStudy (anonymous):

(x^2) (x^2) (x^2) =6

OpenStudy (anonymous):

Just multiply the exponents : 2 * 3 = 6 -> (y^2)^3 = y^6

OpenStudy (anonymous):

Can I continue?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

( x^4 )^3 = ....?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Why 32?

OpenStudy (anonymous):

because u +4 3times

OpenStudy (anonymous):

12 my bad

OpenStudy (anonymous):

Just multiply the exponents : 2 * 3 = 6 -> (y^2)^3 = y^6 -> ( x^4 )^3 = x^12, ( not 12)

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Suppose x is candy: you have 3 candies + 5 candies = ....

OpenStudy (anonymous):

8

OpenStudy (anonymous):

8 candies, ( not 8 )

OpenStudy (anonymous):

Similarly here, pretend x is candies: candies^12